dominic spent 1/4 of his money on a pair of socks, 1/5 of it on a magazine and 50 on a snack. If he had 82 left, how much money did he start with?

Answer by nerdybill(7384)

dominic spent 1/4 of his money on a pair of socks, 1/5 of it on a magazine and 50 on a snack. If he had 82 left, how much money did he start with?
.
Let x = amount he started with
then
x - x/4 - x/5 - 50 = 82
x - x/4 - x/5 = 132
20x/20 - 5x/20 - 4x/20 = 132
(20x-5x-4x)/20 = 132
(20x-9x)/20 = 132
11x/20 = 132
11x = 2640
x = 240